OpenStudy (anonymous):

What is the sum of the arithmetic sequence 4, 11, 18 …, if there are 26 terms?

OpenStudy (anonymous):

Tn=a+(n-1)d d= common difference a= first term n = no. of term Sn=0.5(a+l) l= last term

OpenStudy (anonymous):

d = 7 .thats all i got :(

OpenStudy (anonymous):

d = 7, a = 4, n = 26...

OpenStudy (anonymous):

a=4 d=7 there fore T26=4+(25)7 = 179

OpenStudy (anonymous):

A) 1,482 B) 1,534 C) 2,379 D) 2,470

OpenStudy (anonymous):

those are my choices

OpenStudy (anonymous):

well i did a slight mistake in formula Sn=0.5n(a+l) so Sn=0.5*26*(4+179)

OpenStudy (anonymous):

its 2379

OpenStudy (anonymous):

i got 2327 but thats not a choice

OpenStudy (anonymous):

u r doing wrong its Sn=0.5*26*(4+179)

OpenStudy (anonymous):

u got it now???